OpenStudy (ms_sp):

1. The center circle of a soccer field has a radius of 1 yd. What is the area of the center of circle. Use 3.14. Will medal and fan!

OpenStudy (ms_sp):

@Will.H

OpenStudy (will.h):

Area = pi * r^2 Area = 3.14 * 1^2 Area =?

OpenStudy (ms_sp):

So, why is it not 6 being squared?